Find the area of the circle x2 + y2 = r2.


1
Expert's answer
2021-04-29T17:13:19-0400

Given equation of circle is:

"x^2+y^2=r^2"


"\Rightarrow y=\sqrt{r^2-x^2}"


Area of circle =4 "\times" Area of first quadrant


        "=4 \int_0^rydx"

 

        "=4\int_0^r\sqrt{r^2-x^2}dx"


        "=4[\dfrac{x}{2}\sqrt{r^2-x^2}+\dfrac{r^2}{2}sin^{-1}\dfrac{x}{r}]_0^r"


        "=4[\dfrac{r^2}{2}sin^{-1}-0]"

        

"=4\times \dfrac{r^2\pi}{4}"


        "= \pi r^2 \text{ square units }"